Let f (x) = {√-x ,if x <0
{3-x ,if 0《x <3
{(x-3)^2 ,if x>3
(A) Check whether for is discontinuous.if yes l,find where?
(B) Give a rough sketch of the graph of f.
1
Expert's answer
2020-03-09T11:18:56-0400

"y=\left\{
\begin{matrix}
 \sqrt{-x} , x<0 \\
 3-x, 0\leq x<3\\
(x-3)^2, x>3
\end{matrix}\right."

"x=0\\
\lim\limits_{x\to0-0}\sqrt{-x}=0\\
\lim\limits_{x\to0+0}(3-x)=3\\"

 The function at the point "x=0" has a jump discontinuity


"x=3\\
\lim\limits_{x\to3-0}(3-x)=0\\
\lim\limits_{x\to3+0}(x-3)^2=0\\"

The function  is a continuous at the point "x=3" .

The function  is discontinuous at "x=0".